Present Value of Bonds Payable; Premium Moss Co. issued $280,000 of five-year, 11% bonds, with interest payable semiannually,Exhibit 5 Present Value of $1 at Compound Interest 10% 0.90909 Periods 1 2 3 4 5 6 7 4% 0.96154 0.92456 0.88900 0.85480 0.821 #5

Exhibit 7 Present Value of an Annuity of $1 at Compound Interest Periods 1 2 3 4 4% 0.96154 1.88609 2.77509 3.62990 4.45182 5#7
Answer #1

Present Value of the bonds payable = $ 290,809

Workings:

Selling Price of the bond = Present Value of Lump Sum + Present Value of Interest Payments
Computation of Present Value of Lump Sum
Market Rate of Interest = 5% (10% / 2)
No. of periods = 10 (5 years x 2)
Present Value Factor @ 5% = 0.61391
(a) Present Value of Lump Sum = $       1,71,894.80 ($ 280000 x 0.61391)
Computation of Present Value of Interest Payments
PVAF @ 5% for 10 periods = 7.72173
Interest Expense = $          15,400.00 ($ 280000 x 5.5%)
(b) Present Value of Interest = $       1,18,914.64 ($ 15400 x 7.72173)
(a+b) Issue Price of the bond = $ 171895 + $ 118915
= $             2,90,809

Please let me know if you find this incorrect, will surely help you out. Thank you